OpenStudy (lgbasallote):

Prove by contraposition that if a + b \(\ge\) 15, then a\(\ge\) 8 or b\(\ge\) 8

OpenStudy (aravindg):

8+8=16

OpenStudy (lgbasallote):

...

OpenStudy (lgbasallote):

that's not proving...

OpenStudy (anonymous):

a>=8 .......i b>=8.......ii Adding i and ii a+b>=16

OpenStudy (anonymous):

Since a+b>=15 is always true for a+b>=16. a>=8 and b>=8

OpenStudy (anonymous):

Hmmmmm, you'd have to show that \[a\lt 8 \wedge b \lt 8 \Rightarrow a+b\lt 15\]

OpenStudy (anonymous):

If a < 8 and b < 8 then a + b < 15

OpenStudy (anonymous):

Weird because it doesn't seem true, yet....

ganeshie8 (ganeshie8):

\(a + b \ge 15\) => \(a \ge 8 \) or \(b \ge 8\) is true only if \((a, b) \in \mathbb N \)

OpenStudy (anonymous):

@ganeshie8 What's the counter example, if \(a, b \in \mathbb{R}\)?